DApdr Microaft Word-Axsignment 1.doex QUESTION Q1. An electron, accelerated by a potential difference V, has the de Broglie wavelength 10-10 meter. Calculate the value of that accelerating potential V in volt. Use the standard value of mass of the electron, Planck constant etc wherever needed. (5 Marks) ee to soara
Answer #1

We know that De Broglie wave length for electron is

= h/(2mqv)1/2

m=mass of electron =9.1×10-31 kg

V=accelerating potential

q=charge of electron =1.6×10-19 C

h=Plank's constent 6.62×10-34​​​​​ Joule ×second

So putting the values in above equation

V= h2/2mq×(wavelength)2

  V=(6.62×10​​-34)2/2×9.1×10​-31×1.6×10​​​​​-19×10​​​​​-20

V= 150.495 volt . Answer